1. Let the angles of a triangle be α, β, and γ, with opposite sides of length a, b, and c, respectively. Use the Law of Sines to find the remaining sides. (Round your answers to one decimal place.)

α = 48°;  β = 83°;  c = 112

a=

b=

2. Let the angles of a triangle be α, β, and γ, with opposite sides of length a, b, and c, respectively. Use the Law of Sines to find the remaining sides. (Round your answers to one decimal place.)

β = 105°;  γ = 26°;  c = 24

a=

b=
Answer #1

ab d=40° B=03 c=112 с 2+B+ r = 100 ° r = 100 ° -(x+B) = 100 - (18+83) = 100 - 17 18=497 Law of sine Sin Y Sin 400) a – Sin 49(2) X+B+ 8 = 1000 o = 180°-( 105 +26*) = 180 - 1310 = 49° Law of sine sind a sing ? = Sina > a= Sin & 24 x Sin (49) Sin (26)